# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v3

@@ 4 When will Personalization Improve Statistical Accuracy
The loss function $\ell(\cdot,z)$ is L-smooth, i.e. for any $x,y\in\mathbb{R}^{d}$
$\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\quad\forall z.$ (3)
@@
The loss function $\ell(\cdot,z)$ is almost surely $\mu$-strongly convex: $\ell(x,z)\geq\ell(y,z)+\langle\nabla\ell(y,z),y-x\rangle+\frac{\mu}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{d}.$
@@
There exists a nonnegative constant $\rho$ such that $\forall i\in[m]$, we have $\mathbb{E}_{z\sim\mathcal{D}_{i}}\|\nabla\ell({{\bm{w}}_{\star}^{(i)}},z)\|^{2}\leq\rho^{2}$.
Here, the strongly convex and smooth assumption is used to establish the estimation error and is widely adopted in the theoretical analysis of regularization-based personalized federated learning (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and theoretical analysis in federated learning (Chen et al., 2023b; Cheng et al., 2023). Assumption 4 is used to quantify the observation noise, and is also standard in the literature (Duan & Wang, 2023; Chen et al., 2023b).
Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis. To facilitate the analysis, we first define some notations. Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models $\bm{w}^{(i)}$ in objective (1). A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$. In this paper, we use global statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and local statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as the measure of statstical accuracy. Here the expectation is taken with respect to all data. Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error. The convergence rate of these two definitions is closely related up to the smoothness constant and strong convexity constant. Define $\tilde{\bm{w}}_{\text{GT}}$ as the optimal solution for the GlobalTrain problem in (2). We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error. See Appendix A.2 for the proof.
Lemma 1.
Suppose Assumptions 1, 2 and 3 hold. We have
$\|\widetilde{\bm{w}}^{(g)}-\tilde{\bm{w}}_{\text{GT}}\|^{2}\leq\frac{4\left(1+L/\mu\right)}{\mu(\mu+\lambda)}\sum_{i\in[m]}p_{i}\|\nabla L_{i}(\widetilde{\bm{w}}^{(g)},S_{i})\|^{2},$ (4)
$\sum_{i\in[m],j\in[n_{i}]}\frac{p_{i}}{n_{i}}$ $\left(\mu\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\rangle\right)\leq\frac{\lambda}{2}R^{2}.$ (5)
Lemma 1 establishes a relationship between the model estimation error and the tuning parameter $\lambda$. From the bound (4), we can see that a larger $\lambda$ will push the global model solution $\widetilde{\bm{w}}^{(g)}$ towards the GlobalTrain solution. On the other hand, from the bound (5), we can set a smaller $\lambda\propto 1/nR^{2}$ to obtain a rate $O(1/n)$, independent of $R$. This suggests that $\lambda$ can be chosen adaptively w.r.t. $R$ to achieve the optimal statistical rate. A detailed discussion can be found in Appendix A.3. In the following theorem, we provide such a choice and establish a state-of-the-art statistical rate.
@@
Suppose Assumptions 1, 2, 3 and 4 hold and $n_{i}=n,\forall i\in[m]$. Consider the FedProx objective in (1) with $p_{i}=\frac{1}{m},\forall i\in[m]$. If it takes $\lambda\geq\max\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu$ when $R\leq\frac{1}{\sqrt{n}}$, and $\lambda\leq\frac{\rho^{2}}{n\mu R^{2}}$ when $R>\frac{1}{\sqrt{n}}$, the global model $\widetilde{\bm{w}}^{(g)}$ and local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained from solving the objective (1) satisfy
@@
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\right\|^{2}\leq\frac{C_{1}}{N}+C_{2}\left(\frac{1}{n}\wedge R^{2}\right),$ (6)
@@
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq\frac{C_{3}}{N}+C_{4}(\frac{1}{n}\wedge R^{2})\quad\text{ forall }i\in[m],$ (7)
where $N=mn$ and
$C_{1}=32\frac{\rho^{2}}{\mu^{2}},C_{2}=\big{(}\frac{(1+\sqrt{3})\rho}{\mu}\big{)}^{2}\vee\frac{32L^{2}}{\mu^{2}},C_{3}=\frac{192\rho^{2}}{\mu^{2}},C_{4}=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee\frac{198L^{2}}{\mu^{2}}.$
Proof.
See Appendix A.3. ∎
Theorem 1 shows that with an appropriate choice of $\lambda$, both the global and local models achieve a fast statistical rate across a wide range of $R$. To interpret the result, we first analyze the error bound provided in (6) and (7). The two error bounds have the same order and can be decomposed into two key components. The first term, of order $\mathcal{O}(N^{-1})$, captures the benefit of collaborative learning, as $N=mn$ is the total sample size across $m$ clients. This term reflects the sample efficiency gained through collaborative learning. The second term, of order $\mathcal{O}(n^{-1}\wedge R^{2})$, measures the cost of statistical heterogeneity. As $R$ increases, the convergence rate gradually degrades, transitioning from $\mathcal{O}(N^{-1})$ to $\mathcal{O}(n^{-1})$. Notably, the guarantee is minimax-optimal, as the derived bound matches the problem’s lower bound under the given assumptions (Chen et al., 2023b). Remark 2 in Appendix A.3 discusses about how such an interpolation is achieved through different personalization degrees.
Table 1: Results Comparison. $D:=\sup_{\bm{w},\bm{w}^{\prime}\in\mathcal{D}}\|\bm{w}-\bm{w}^{\prime}\|$ is diameter of the parameter space ($D\geq R$) and $\|\ell\|_{\infty}=\inf\{M\in\mathbb{R}:\ell(\cdot,z)\leq M,\text{for any }z\}$ is an uniform upper bound on the loss function. Statistical rate refers to $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, with constants neglected.
Source Assumption Paradigm Statistical Rate
Chen et al. (2023b) A1,2,3,4, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
Ours A1,2,3,4 One-stage $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
Lower Bound - - $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
We now compare our theoretical results with the result up to date (Chen et al., 2023b), as summarized in Table 1. For a fair comparison, it is noticeable that we have transformed their result into our metrics $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, by removing their optimization error and problem-dependent constant. Our contribution can be summarized into three aspects:
$\bullet$ Our results are derived under more realistic assumptions. Chen et al. (2023b) imposes additional stringent conditions including a bounded parameter space $D$ and a uniform bound on the loss function $\ell(\cdot,z)$. Such conditions are difficult to satisfy in practice. For example, even a simple linear regression model with sub-Gaussian covariates would violate these assumptions.
$\bullet$ Our result is established for both global and local models with the same choice of $\lambda$. In contrast, Chen et al. (2023b) proposes a two-stage process where the global model is first estimated, followed by an additional local training phase with a different choice of $\lambda$. This two-step approach is likely the artifact of their analysis, as existing work on FedProx also shows that a single-stage optimization with a chosen $\lambda$ suffices to achieve the desired performance (T Dinh et al., 2020).
$\bullet$ We establish a state-of-the-art minimax statistical rate. As shown in Table 1, the established upper bound matches the lower bound established in Chen et al. (2023b). In a heterogeneous case when $R\geq n^{-1/2}$, Chen et al. (2023b) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$. This bound becomes increasingly loose with larger $R$. In contrast, our bound is $\mathcal{O}\left(n^{-1}\right)$, which shows that properly-tuned FedProx is always no worse than LocalTrain, independent of $R$. In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data. As the client number could be extremely large ($10^{5}$ devices in chen2023fs), order of $m$ is non-trivial. During the transition period when $m^{-1}n^{-1/2}<R<n^{-1/2}$, we achieve a rate of $\mathcal{O}(R^{2})$, again matching the lower bound, and is strictly faster than the rate $\mathcal{O}(n^{-1/2}R)$ established in Chen et al. (2023b).
5 When will Personalization Improve Communication Efficiency
@@
Notice that if we define the local objective of each client $i$ as $h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}):={L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}$, then the FedProx problem given by (1) can be rewritten in the following bilevel (iterated minimization) form:
$\min_{{\bm{w}}^{(g)}}F({\bm{w}}^{(g)}):=\frac{1}{m}\sum_{i=1}^{m}F_{i}({\bm{w}}^{(g)}),\quad\text{where}\quad F_{i}({\bm{w}}^{(g)}):=\min_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (8)
@@
The reformulation (8) is a finite-sum minimization problem, with each component $F_{i}$ being the Moreau envelope (Moreau, 1965; Yosida, 1964) of $L_{i}$. Let ${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})$ be the minimizer of $h_{i}(\,\cdot\,,{\bm{w}}^{(g)})$, i.e.,
${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)}):=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})=\operatorname*{arg\,min}_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (9)
Based on Assumption 2 and the new notations in (8) and (9), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in Appendix B.2.
@@
Next, we recall the properties of the Moreau envelope to guide the algorithm design.
@@
From Lemma 2 and Lemma 3, it suggests applying a simple gradient algorithm to optimize ${\bm{w}}^{(g)}$:
$\begin{split}{\bm{w}}_{t+1}^{(g)}&={\bm{w}}_{t}^{(g)}-\gamma\cdot\nabla F({\bm{w}}_{t}^{(g)})={\bm{w}}_{t}^{(g)}-\gamma\lambda\cdot\frac{1}{m}\sum_{i=1}^{m}\left({\bm{w}}_{t}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})\right),\end{split}$ (10)
where $\bm{w}_{t}^{(g)}$ is the update at the communication round $t$, $\gamma>0$ is a step size to be properly set (cf. Theorem 2). To implement (10), in each communication round $t$ the server broadcasts ${\bm{w}}_{t}^{(g)}$ to all clients, then each client $i$ updates its local model ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ and upload to the server.
Note that executing (10) requires each client $i$ computing the minimizer ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$. In general, the subproblem (9) does not have a closed-form solution. Therefore, computing the exact gradient $\nabla F$ will incur a high computation cost as well as high latency for the learning process. Leveraging recent advancements in bilevel optimization algorithm design (Ji et al., 2022), we address this issue by approximating ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ with a finite number of $K$ gradient steps. Specifically, we let each client $i$ maintain a local model ${\bm{w}}_{t,k}^{(i)}$. Per communication round $t$, ${\bm{w}}_{t,k}^{(i)}$ is initialized to be ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$ as a warm start, and updated according to
$\begin{split}{\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\eta\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)}),\quad\forall\,k=0,\ldots,K-1,\end{split}$ (11)
where $\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)})$, for notation simplicity, denotes the partial gradient of $h_{i}$ with respect to ${\bm{w}}^{(i)}$. The overall procedure is summarized in Alg. 1, see Appendix B.1.
Remark 1.
Alg. 1 is a customization of the general bi-level problem (Ji et al., 2022). While the algorithm design is not new, our contribution lies in analyzing how the personalization degree influences the trade-off between communication and local computation leveraging the FedProx problem structure. Extending to noisy settings to consider stochastic gradient or variance reduction would be interesting future work, which requires additional technical effort.
@@
Proof.
See Appendix B.3 and B.5. ∎
Theorem 2 shows under the step size constraint $\gamma<1/L_{g}$, ${\bm{w}}_{t}^{(g)}$ converges to $\tilde{{\bm{w}}}^{(g)}$ at the same rate as the iteration (10) with exact gradient computation. Furthermore, the larger $\gamma$ is, the larger $K$ should be chosen so as to satisfy (12). This indicates a tradeoff between communication and computation: the fewer communication rounds are performed, then the more local computation is required for the algorithm to converge.
To better interpret the result and obtain insights on the influence of $\lambda$ on the convergence, we provide the communication and computation complexity under specific choices the algorithm’s tuning parameters, as described in the following Corollary 1. See Appendix B.4 for the proof.
@@
In the same setting as Theorem 2, if choosing the step size $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and the inner loop iteration number $K=\tilde{\mathcal{O}}\left(\frac{\lambda+L}{\lambda+\mu}\right)$, then the communication complexity and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\tilde{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\tilde{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
@@
(i) # communication $\begin{split}&\text{(i) \# communication rounds}=\mathcal{O}\big{(}\kappa_{g}\log(1/\varepsilon)\big{)}=\mathcal{O}\left(\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right),\end{split}$ (13)

### v4

@@ 4.1 Effect of Personalization on Statistical Accuracy
The loss function $\ell(\cdot,z)$ is L-smooth, i.e. for any $x,y\in\mathbb{R}^{d}$:
$\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\ \forall z.$ (6)
@@
The empirical loss $L_{i}(\cdot,S_{i})$ is $\mu$-strongly convex, i.e., for any $x,y\in\mathbb{R}^{d}$ and $i\in[m]$:
$L_{i}(x,S_{i})\geq L_{i}(y,S_{i})+\langle\nabla L_{i}(y,S_{i}),x-y\rangle+\frac{\mu}{2}\|x-y\|^{2}.$ (7)
@@
There exists a nonnegative constant $\rho$ such that $\mathbb{E}_{z\sim\mathcal{D}_{i}}\|\nabla\ell({{\bm{w}}_{\star}^{(i)}},z)\|^{2}\leq\rho^{2}$ for all $i\in[m]$.
The strong convexity and smoothness assumptions are used to establish the estimation error and are widely adopted in the theoretical analysis of regularization-based PFL (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and FL (Chen et al., 2023c; Cheng et al., 2023). Assumption 3 is used to quantify the observation noise, and is also standard in the literature (Duan & Wang, 2023; Chen et al., 2023c).
For simplicity, we assume $p_{i}=1/m$ and $n_{i}=n$ for all $i\in[m]$. The following theorem provides an upperbound of the estimation error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as a function of $\lambda$.
@@
Suppose Assumption 1, 2 and 3 hold and consider the parameter space $\mathcal{P}(R)$ given by (5). The local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained by solving Problem (2) satisfy
@@
$\begin{split}&\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &\leq\min\Bigg{\{}\frac{48\rho^{2}}{\mu^{2}}\frac{1}{N}+\left(\frac{48L^{2}}{\mu^{2}}+3\right)R^{2}+\frac{1}{q_{1}(\lambda)},\\ &\frac{4\rho^{2}}{\mu^{2}}\frac{1}{n}+\left[\left(\frac{4\rho^{2}}{\mu^{3}}\frac{1}{n}+\frac{4}{\mu}R^{2}\right)\lambda+\frac{8}{\mu^{2}}R^{2}\lambda^{2}\right]\Bigg{\}},\end{split}$ (8)
where $N=mn$ and $q_{1}(\lambda)$, defined in (70), is a monotonically increasing function of $\lambda$ with $\lim_{\lambda\to\infty}q_{1}(\lambda)=\infty$.
See Appendix A.3 for the proof. The bound given by (8) consists of two terms. The first term decreases as $\lambda$ increases. When $\lambda\to\infty$, the term $[q_{1}(\lambda)]^{-1}\to 0$ and the first term approaches $\mathcal{O}(1/N+R^{2})$. Here, $\mathcal{O}(1/N)$ corresponds to the sample complexity one could obtain if the data distributions of the $m$ clients are homogeneous, reflecting the benefit of collaborative training. Term $\mathcal{O}(R^{2})$ reflects the negative impact due to the bias introduced by statistical heterogeneity. In contrast, the second term in (8) increases with $\lambda$, showing that a smaller $\lambda$ leads to a rate closer to $\mathcal{O}(1/n)$ corresponding to that of pure local training. This rate is independent of $R$, making it robust to high data heterogeneity but at the expense of reduced sample efficiency. Together, these results provide a continuous and quantitative characterization of how the personalization degree, governed by $\lambda$, determines the statistical accuracy of the solution $\widetilde{\bm{w}}^{(i)}$.
Minimax-optimal Statistical Accuracy. To demonstrate the tightness of our analysis, we now show that as a direct implication of Theorem 1, the local models $\widetilde{\bm{w}}^{(i)}$ are rate-optimal. Denoting $\kappa=L/\mu$ as the conditional number, we have the following corollary.
Corollary 1.
Suppose Assumption 1, 2 and 3 hold and consider the parameter space $\mathcal{P}(R)$ given by (5). If setting $\lambda\geq\max\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu$ when $R\leq\frac{1}{\sqrt{n}}$, and $\lambda\leq\frac{\rho^{2}}{n\mu R^{2}}$ when $R>\frac{1}{\sqrt{n}}$, then for all $i\in[m]$, the local model $\widetilde{\bm{w}}^{(i)}$ obtained by solving Problem (2) satisfy
@@
$\begin{split}\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq C_{3}\frac{1}{N}+C_{4}\left(R^{2}\wedge\frac{1}{n}\right),\quad\end{split}$ (9)
where constants $C_{3}$ and $C_{4}$ are defined in (91).
See Appendix A.3 for the proof. In addition, Theorem 8 in Chen et al. (2023c) states for all $i\in[m]$ and any estimator $\hat{\bm{w}}^{(i)}$ that is a measurable function of data $\{S_{i}\}_{i=1}^{m}$, we have
$\begin{split}&\inf_{\hat{\bm{w}}^{(i)}}\sup_{\{{{\bm{w}}_{\star}^{(i)}}\}_{i=1}^{m}\in\mathcal{P}(R)}\mathbb{E}\left\|\hat{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\gtrsim\frac{1}{N}+R^{2}\wedge\frac{1}{n}.\end{split}$
Therefore, Corollary 1 shows that with an appropriate choice of $\lambda$, the established rate achieves the minimax lower bound. A detailed comparison with prior results can be found in Appendix A.5, along with a discussion of several novel techniques developed to derive our result. While this result focuses on statistical accuracy and is not directly tied to the accuracy-communication trade-off, it highlights the tightness of our analysis. Furthermore, the techniques introduced to obtain the results are of independent interest. To the best of our knowledge, this is the first work to establish minimax-optimality for the solutions of Problem (2).
Remark 2 (Technical Novelty of The Analysis).
Unlike previous results (Chen et al., 2023c) which are algorithm-dependent, our analysis does not rely on any algorithm but directly tackles the objective function, establishing rates using purely the properties of the loss. Specifically, using the strong convexity of the loss function, we first proved the estimation error is bounded by the gradient norm and an extra statistical heterogeneity term controlled by $\lambda$, as detailed in Equation (33). This allows us to show for large $R$, a small $\lambda$ can be chosen to yield rate $O(1/n)$ and match one of the worst cases in the lower bound. For the complementary case $R\leq 1/\sqrt{n}$, we leveraged the GlobalTrain solution $\widetilde{\bm{w}}_{GT}$ as a bridge and proved the solutions of Problem (2) to $\widetilde{\bm{w}}_{GT}$ are bounded by a term inversely proportional to $\lambda$ (cf. Lemma 4), implying that they can be made arbitrarily close to $\widetilde{\bm{w}}_{GT}$ by setting $\lambda$ small. This, together with the rate of $\widetilde{\bm{w}}_{GT}$, yields a rate of $O(1/N+R^{2})$. Combining the two cases, we proved minimax optimality. Not only each piece of the above results are new, but more importantly, identifying they are the key ingredients to show the solution of (2) is minimax optimal, are the technical novelties of this proof.
Notice that Theorem 1 and Corollary 1 focus on establishing the statistical rate for the solution in Problem (2), $\widetilde{\bm{w}}^{(i)}$. To fully quantify the error of the output of PFL algorithms, we need to establish the optimization error from the algorithm as well, as detailed in the next section.
4.2 Effect of Personalization on Communication Efficiency
In this section, we provide an FL algorithm for solving Problem (2) along with its convergence analysis, showing the impact of the personalization degree on communication and computation complexity. Notice that if we define the local objective of each client $i$ as
$h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}):={L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2},$
@@ 4.2 Effect of Personalization on Communication Efficiency
then Problem (2) can be rewritten in the following bilevel (iterated minimization) form:
$\begin{split}&\min_{{\bm{w}}^{(g)}}F({\bm{w}}^{(g)}):=\frac{1}{m}\sum_{i=1}^{m}F_{i}({\bm{w}}^{(g)}),\\ &\text{where}\quad F_{i}({\bm{w}}^{(g)}):=\min_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).\end{split}$ (10)
@@
The reformulation (10) has a finite-sum minimization structure, with each component $F_{i}$ being the Moreau envelope (Moreau, 1965; Yosida, 1964) of $L_{i}$. Let ${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})$ be the minimizer of $h_{i}(\,\cdot\,,{\bm{w}}^{(g)})$, i.e.,
$\begin{split}{\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})&=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})\\ &:=\operatorname*{arg\,min}_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).\end{split}$ (11)
The following lemmas provide properties of $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$, instrumental to the algorithm design and rate analysis. The proof can be found in Appendix B.2.
@@
Lemma 1 and Lemma 2 suggest applying a simple gradient algorithm to optimize ${\bm{w}}^{(g)}$:
$\begin{split}{\bm{w}}_{t+1}^{(g)}&={\bm{w}}_{t}^{(g)}-\gamma\cdot\nabla F({\bm{w}}_{t}^{(g)})\\ &={\bm{w}}_{t}^{(g)}-\gamma\lambda\cdot\frac{1}{m}\sum_{i=1}^{m}\left({\bm{w}}_{t}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})\right),\end{split}$ (12)
where $\bm{w}_{t}^{(g)}$ is the update at the communication round $t$, $\gamma>0$ is a step size to be properly set (cf. Theorem 2). To implement (12), in each communication round $t$ the server broadcasts ${\bm{w}}_{t}^{(g)}$ to all clients, then each client $i$ updates its local model ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ and uploads to the server.
Note that executing the update (12) requires each client $i$ computing the minimizer ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$. In general, the subproblem (11) does not have a closed-form solution. Therefore, computing the exact gradient $\nabla F$ will incur a high computation cost as well as high latency for the learning process. Leveraging recent advancements in bilevel optimization algorithm design (Ji et al., 2022), we address this issue by approximating ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ with a finite number of $K$ gradient steps. Specifically, we let each client $i$ maintain a local model ${\bm{w}}_{t,k}^{(i)}$. Per communication round $t$, ${\bm{w}}_{t,k}^{(i)}$ is initialized to be ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$ as a warm start, and updated according to
$\begin{split}{\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\eta\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)}),\end{split}$ (13)
for $k=0,\ldots,K-1$, where $\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)})$, for notation simplicity, denotes the partial gradient of $h_{i}$ with respect to ${\bm{w}}^{(i)}$. The overall procedure is summarized in Alg. 1, see Appendix B.1.
@@
The proof of Theorem 2 can be found in Appendix B.3 and B.3.2. Theorem 2 shows that with sufficiently small step sizes, both the local and global models converge linearly. The personalization degree, controlled by $\lambda$, plays a critical role in influencing the algorithm’s communication efficiency. A small $\lambda$ (indicating higher personalization) yields a small $L_{g}$, and thus permits a larger choice of the step size $\gamma$. By the expression of the convergence rate $21-(\gamma\mu_{g})/2-(\gamma\mu_{g})^{2}/$ 2, we can see less communication round would be required to reach an $\varepsilon$-optimal solution. However, increasing $\gamma$ will also decrease the right hand side of (14), implying that a larger number of local gradient steps $K$ should be taken to fulfill the condition. To provide a more concrete characterization of these dynamics, we present Corollary 2, which quantifies the communication and computation complexities under specific choices of tuning parameters. The proof of Corollary 2 can be found in Appendix B.4.
@@
In the setting of Theorem 2, if we further choose the step size $\eta=(\lambda+L)^{-1}$, $\gamma=(\lambda+L)/(2\lambda L)$, and the inner loop iteration number $K=\tilde{\mathcal{O}}((\lambda+L)/(\lambda+\mu))$, then ${\bm{w}}_{t}^{(g)}$ converges to $\widetilde{\bm{w}}^{(g)}$ linearly at rate $1-(\lambda+L)/(4\kappa(\lambda+\mu))$, and ${\bm{w}}_{t,0}^{(i)}$ converges to $\widetilde{\bm{w}}^{(i)}$ linearly at the same rate for any $i\in[m]$. Thus the communication and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\tilde{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\tilde{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
@@
(i) # communication $\begin{split}&\text{(i) \# communication rounds}=\mathcal{O}\Big{(}\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\Big{)},\end{split}$